The ratio of male students to female students in the drama club at Campbell High School is 3:4. If the number of male students in the club is 18,predict the number of female students?
Answer by jim_thompson5910(35256)

3/4 = 18/x

3x = 4*18

3x = 72

x = 72/3

x = 24


So there are 24 female students.
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------